# On the Duality between Gradient Transformations and Adapters

### v1

@@ 5 Discussion and Limitations
The preceding studies focus on two situations in which the duality between linear adapters and gradient transformations offers practical insights. We believe there are many other avenues that merit further exploration. For instance, § 3.1 makes no assumptions about the structure of ${\mathbf{S}}$; while we only considered Kronecker-factorized matrices, other linear maps that admit efficient storage and computation would be interesting to explore. Regardless of the structure of ${\mathbf{S}}$, as discussed in § 4.1, what characterizes a good ${\mathbf{S}}$ is not clear but has a large impact. It may be possible to learn a good ${\mathbf{S}}$ with meta-learning-style approaches, which can be seen as learning an optimizer (Andrychowicz et al., 2016; Li & Malik, 2016; Wichrowska et al., 2017; Bello et al., 2017, i.a.). [^12] In the GaLore/LoRA case, learning ${\mathbf{P}}$ in this meta-learning sense is different from learning ${\mathbf{P}}$ in the ordinary LoRA sense, i.e., when both ${\mathbf{P}}$ and ${\mathbf{A}}$ are trained with gradient descent against the same loss function. Finally, while we focused on linear gradient transformations, where we proved exact equivalence with a linear adapter parameterization, it may be possible to establish approximate equivalence between non-linear gradient transformations and other types of adapters.
@@ Acknowledgements
We thank Shannon Zejiang Shen, Li Du, Aniruddha Nrusimha, Jeremy Bernstein, Jyothish Pari, Sami Jaghouar, and Johannes Hagemann for helpful discussions and feedback. This study was supported by MIT-IBM Watson AI Lab.
@@ Appendix A Proofs
where we used eq. 3 in the first equality, and eq. 4 for the second equality. Also by eq. 4 and by the chain rule, we have that the gradients of the loss function at timestep $t$ to the (effective) parameters of the networks are the same, i.e., $\overline{\Theta^{(0)}+{\mathbf{S}}^{\top}\Lambda^{(k)}}=\overline{\Theta^{(k)}}.$ In particular, this means that
@@
and in turn, expanding $\Delta_{\Lambda}^{(k)}$,
@@
where the first equality is the optimizer we use to train the reparameterized model in eq. 3, in the second equality we use eq. 7 and eq. 5. But by eq. 6,
@@
Eq. 8 proves optimizer states are equivalent for optimizer step $k+1$, whereas eq. 9 establishes the networks are equivalent for optimizer step $k+1$, thus completing the proof. ∎
A.2 Proof of § 3.2
@@
From § 3.1, we know that training a linear layers using the gradient transformation ${\mathbf{S}}={\mathbf{R}}^{\top}\otimes{\mathbf{L}}$ corresponds to using the reparameterization:
@@
A.3 Proof of § 3.2
We now state a more general version of § 3.2, and then prove it.
@@
$\widetilde{\overline{{\mathbf{W}}}}=\begin{cases}\mathbf{P}\overline{{\mathbf{W}}}&m\leq n\,\,\,\,\,\,\text{(i.e., apply from the left)}\\ \overline{{\mathbf{W}}}{\mathbf{P}}&m>n\,\,\,\,\,\,\text{(i.e., apply from theright)}\\ \end{cases}$
@@
${\mathbf{W}}^{(t+1)}=$$\begin{cases}{\mathbf{W}}^{(t)}+{\mathbf{P}}^{\top}\operatorname{vec}^{-1}\left(\Delta_{{\mathbf{W}}}^{(t)}\right)&m\leq n\\ {\mathbf{W}}^{(t)}+\operatorname{vec}^{-1}\left(\Delta_{{\mathbf{W}}}^{(t)}\right){\mathbf{P}}^{\top}&m>n\\ \end{cases}$
@@
${\mathbf{S}}=\begin{cases}{\mathbf{I}}_{n}\otimes{\mathbf{P}}&m\leq n\\ {\mathbf{P}}^{\top}\otimes{\mathbf{I}}_{m}&m>n\\ \end{cases}$
@@
$\operatorname{vec}\left(\widetilde{\overline{{\mathbf{W}}}}\right)$ =$=\begin{cases}\operatorname{vec}\left(\mathbf{P}\overline{{\mathbf{W}}}\right)&m\leq n\\ \operatorname{vec}\left(\overline{{\mathbf{W}}}{\mathbf{P}}\right)&m>n\\ \end{cases}$
@@
using $\operatorname{vec}\left({\mathbf{M}}{\mathbf{N}}{\mathbf{O}}\right)=({\mathbf{O}}^{\top}\otimes{\mathbf{M}})\operatorname{vec}\left({\mathbf{N}}\right)$ as before. imilarly, we have that $\operatorname{vec}\left({\mathbf{W}}^{(t+1)}\right)=\operatorname{vec}\left({\mathbf{W}}^{(t)}\right)+{\mathbf{S}}^{\top}\Delta_{{\mathbf{W}}}^{(t)}$. Hence, by § 3.1, we have that training a network with GaLore is equivalent to introducing a parameter ${\mathbf{A}}$ and optimizing using the reparameterization $\operatorname{vec}\left({\mathbf{W}}\right)=\operatorname{vec}\left({\mathbf{W}}^{(0)}\right)+{\mathbf{S}}\operatorname{vec}\left({\mathbf{A}}\right)$. Observe that this choice of ${\mathbf{S}}$ is a special case of § 3.2 where ${\mathbf{L}}={\mathbf{P}}$ and ${\mathbf{R}}={\mathbf{I}}$ (if $m\leq n$) or ${\mathbf{L}}={\mathbf{I}}$ and ${\mathbf{R}}={\mathbf{P}}$ (if $m>n$). Thus, the reparameterization corresponds to LoRA with one of the two adapter matrices frozen to ${\mathbf{P}}$. ∎
@@ Appendix B Weight Decay
§ 3.2 establishes an equivalence between GaLore and LoRA when stateful optimizers are in play (eq. 2). While Adam (Kingma & Ba, 2015) can be straightforwardly recast as a stateful optimizer, it turns out that weight decay, as is traditionally implemented in, e.g., AdamW (Loshchilov & Hutter, 2019), breaks this symmetry as it does not fit our definition of a stateful optimizer. From the perspective of our definition, the problem is that optimizer steps with AdamW are not solely a function of the observed gradients up until this point, but also the actual values of the parameters. This is important, since automatic differentiation libraries traditionally distinguish trainable and non-trainable parameters, with weight decay being applied to the former. Since the duality in § 3.2 changes what the trainable parameters are, this means that the weight decay is applied differently in the gradient transformation view and in the linear adapter view.
@@
[…]
@@
where $\lambda\in\mathbb{R}^{+}$ is our weight decay penalty. Similarly, following § 3.1, our linear layer’s effective weight after a single optimizer step, in the adapter view, is given by
@@
where the final equality follows from § 3.1. Note that in general, we do not have that $\Theta^{(0)}={\mathbf{S}}^{\top}\Lambda^{(0)}$, which means the optimizer trajectories may diverge.
@@
If one truly cares about preserving the optimizer trajectory even when training with weight decay, practically all one has to do is adjust the application of the weight decay so that it reflects the behavior of weight decay in the gradient transformation view or in the linear adapter view. Adjusting the linear adapter weight decay application to match the gradient transformation weight decay application is fairly simple: one just has to compute the effective weight at every timestep, as done above, and decay the frozen base weights directly. The converse is possible but slightly trickier, since the application of weight decay in the linear adapter view requires one to know what $\Lambda^{(t)}$ is, which may require the introduction of additional optimizer state. For example, one could store $\Theta^{(0)}$ and solve [^13] Note that since the right hand side lies in $\operatorname{range}({\mathbf{S}}^{\top})$, this linear system will have a solution. ${\mathbf{S}}^{\top}\Lambda_{\text{effective}}^{(t)}=\Theta^{(t)}-\Theta^{(0)}$ on every optimizer state,141414 In the distributed case, this might not incur any additional cost, since $\Theta^{(0)}$ would already need to be stored. But this would require solving a linear system on every iteration. or one could store and continually update $\Lambda_{\text{effective}}$ as part of the optimizer state.
Appendix C Experimental Setup
The details of the two architectures we consider are shown in Tab. 5. We include discussion on some additional details below.
@@ Appendix C Experimental Setup
The experimental setup in the original GaLore paper did not perform gradient accumulation, which meant that the maximum sequence length had to be short enough (e.g., 256) such that a single batch could contain a large-enough number of sequences for accurate gradient estimation. Our experiments are instead conducted in the standard setting where we assume gradients are accumulated across multiple microbatches. In this case, the reparameterization of GaLore as LoRA has the additional benefit of straightforwardly allowing for gradient accumulation in the lower-dimensional space. Concretely, the most straightforward implementation of GaLore [^15] E.g., the official implementation in https://github.com/jiaweizzhao/GaLore. will lead to gradient accumulation in the original parameter space, which would consume substantially more memory. In contrast, in the LoRA formulation the gradients of ${\mathbf{A}}$ are accumulated after applying the gradient transformation, providing substantial memory savings without any additional code. [^16] One can implement this optimization in the GaLore form, but this requires additional code. The issue is that most deep learning frameworks will compute the gradients of a parameterized function, and the user then separately passes these as input to an optimizer. If gradients are only transformed in the optimizer, then the automatic differentiation module cannot figure out that only the smaller transformed gradient is needed, and not the full gradient. We suspect that a sufficiently good compiler should in principle recover this optimization if one ensures that entire training steps (viz., all the gradient accumulation steps and optimizer step) are compiled jointly.
@@
Appendix D Additional results
Figure 2: Full results for gradient reconstruction error (i.e., $\|\overline{\Theta}-{\mathbf{S}}^{\top}{\mathbf{S}}\overline{\Theta}\|^{2}$) (left) and cosine similarity (i.e., $\cos(\overline{\Theta},{\mathbf{S}}^{\top}{\mathbf{S}}\overline{\Theta})$) of the various transformations across training steps with the 200M model. The projections with lowest reconstruction error (measured either by L2 error or cosine similarity with the unprojected stochastic gradient) do not give the best downstream performance (see Tab. 2).

### v2

@@ 5 Discussion and Limitations
The preceding studies focus on two situations in which the duality between linear adapters and gradient transformations offers practical insights. We believe there are many other avenues that merit further exploration. For instance, §˜3.1 makes no assumptions about the structure of $\mathbf{S}$}; while we only considered Kronecker-factorized matrices, other linear maps that admit efficient storage and computation would be interesting to explore. The transformation also just needs to be applied to a gradient vector, suggesting that one could compress the gradient of multiple layers in a network jointly. This amounts to sharing LoRA parameters across layers on the GaLore–LoRA case, which has been found to be successful for parameter-efficient finetuning (Renduchintala et al., 2024; Song et al., 2024). eegardless of the structure of $\mathbf{S}$}, as discussed in §˜4.1, what characterizes a good $\mathbf{S}$} is not clear but has a large impact. It may be possible to learn a good $\mathbf{S}$} with meta-learning-style approaches, which can be seen as learning an optimizer (Andrychowicz et al., 2016; Li & Malik, 2016; Wichrowska et al., 2017; Bello et al., 2017, i.a.). [^14] In the GaLore/LoRA case, learning $\mathbf{P}$} in this meta-learning sense is different from learning $\mathbf{P}$} in the ordinary LoRA sense, i.e., when both $\mathbf{P}$} and $\mathbf{A}$} are trained with gradient descent against the same loss function. Finally, while we focused on linear gradient transformations, where we proved exact equivalence with a linear adapter parameterization, it may be possible to establish approximate equivalence between non-linear gradient transformations and other types of adapters.
@@ Acknowledgements
We thank Shannon Zejiang Shen, Li Du, Aniruddha Nrusimha, Jeremy Bernstein, Jyothish Pari, Sami Jaghouar, Johannes Hagemann, and the anonymous reviewers for helpful discussions and feedback. This work was partially supported by the National Science Foundation under CAREER Award No. 2441872 and the MIT-IBM Watson AI Lab.
@@ Appendix A Proofs
where we used eq.˜3 in the first equality, and eq.˜4 for the second equality. Also by eq.˜4 and by the chain rule, we have that the gradients of the loss function at timestep $t$ to the (effective) parameters of the networks are the same, i.e., $\overline{\Theta^{(0)}+{\mathbf{S}}^{\top}\Lambda^{(k)}}=\overline{\Theta^{(k)}}.$ In particular, this means that
@@
where the first two equalities follow from the chain rule. Specifically, the first equality follows from writing the derivative of the loss w.r.t. the adapter parameter $\Lambda^{(k)}$ as a function of the derivative of the loss w.r.t. the additive component of the adapter, ${\mathbf{S}}^{\top}\Lambda^{(k)}$; the second equality follows from recognizing that the latter is equal to the derivative of the loss w.r.t. the effective value $\Theta^{(0)}+{\mathbf{S}}^{\top}\Lambda^{(k)}$ of the (adapted) parameter. In turn, expanding $\Delta_{\Lambda}^{(k)}$,
@@
where the first equality is the optimizer we use to train the reparameterized model in eq.˜3, in the second equality we use eq.˜7 and eq.˜5. But by eq.˜6,
@@
Eq.˜8 proves optimizer states are equivalent for optimizer step $k+1$, whereas eq.˜9 establishes the networks are equivalent for optimizer step $k+1$, thus completing the proof. ∎
A.2 Proof of §˜3.2
@@
From §˜3.1, we know that training a linear layers using the gradient transformation ${\mathbf{S}}={\mathbf{R}}^{\top}\otimes{\mathbf{L}}$ corresponds to using the reparameterization:
@@
A.3 Proof of §˜3.2
We now state a more general version of §˜3.2, and then prove it.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
using $\operatorname{vec}\left({\mathbf{M}}{\mathbf{N}}{\mathbf{O}}\right)=({\mathbf{O}}^{\top}\otimes{\mathbf{M}})\operatorname{vec}\left({\mathbf{N}}\right)$ as before. Similarly, we have that $\operatorname{vec}\left({\mathbf{W}}^{(t+1)}\right)=\operatorname{vec}\left({\mathbf{W}}^{(t)}\right)+{\mathbf{S}}^{\top}\Delta_{{\mathbf{W}}}^{(t)}$. Hence, by §˜3.1, we have that training a network with GaLore is equivalent to introducing a parameter $\mathbf{A}$} and optimizing using the reparameterization $\operatorname{vec}\left({\mathbf{W}}\right)=\operatorname{vec}\left({\mathbf{W}}^{(0)}\right)+{\mathbf{S}}\operatorname{vec}\left({\mathbf{A}}\right)$. Observe that this choice of $\mathbf{S}$} is a special case of §˜3.2 where ${\mathbf{L}}={\mathbf{P}}$ and ${\mathbf{R}}={\mathbf{I}}$ (if $m\leq n$) or ${\mathbf{L}}={\mathbf{I}}$ and ${\mathbf{R}}={\mathbf{P}}$ (if $m>n$). Thus, the reparameterization corresponds to LoRA with one of the two adapter matrices frozen to $\mathbf{P}$}. ∎
@@ Appendix B Weight Decay
§˜3.2 establishes an equivalence between GaLore and LoRA when stateful optimizers are in play (i.e., those satisfying eq.˜2). While Adam (Kingma & Ba, 2015) can be straightforwardly recast as a stateful optimizer, it turns out that weight decay, as is traditionally implemented in, e.g., AdamW (Loshchilov & Hutter, 2019), breaks this symmetry as it does not fit our definition of a stateful optimizer. From the perspective of our definition, the problem is that optimizer steps with AdamW are not solely a function of the observed gradients up until this point, but also the actual values of the parameters. This is important, since automatic differentiation libraries traditionally distinguish trainable and non-trainable parameters, with weight decay being applied to the former. Since the duality in §˜3.2 changes what the trainable parameters are, this means that the weight decay is applied differently in the gradient transformation view and in the linear adapter view.
@@
$(\Delta_{\Theta}^{(0)},\xi_{\Theta}^{(1)})$ $=\text{OptimizerWithoutWeightDecay}({\mathbf{S}}\overline{\Theta}^{(0)},\xi_{\Theta}^{(0)})$ (10)
@@
where $\lambda\in\mathbb{R}^{+}$ is our weight decay penalty. Similarly, following §˜3.1, our linear layer’s effective weight after a single optimizer step, in the adapter view, is given by
@@
where the final equality follows from §˜3.1. Note that in general, we do not have that $\Theta^{(0)}={\mathbf{S}}^{\top}\Lambda^{(0)}$, which means the optimizer trajectories may diverge.
@@
If one truly cares about preserving the optimizer trajectory even when training with weight decay, practically all one has to do is adjust the application of the weight decay so that it reflects the behavior of weight decay in the gradient transformation view or in the linear adapter view. Adjusting the linear adapter weight decay application to match the gradient transformation weight decay application is fairly simple: one just has to compute the effective weight at every timestep, as done above, and decay the frozen base weights directly. The converse is possible but slightly trickier, since the application of weight decay in the linear adapter view requires one to know what $\Lambda^{(t)}$ is, which may require the introduction of additional optimizer state. For example, one could store $\Theta^{(0)}$ and solve [^15] Note that since the right hand side lies in $\operatorname{range}({\mathbf{S}}^{\top})$, this linear system will have a solution. ${\mathbf{S}}^{\top}\Lambda_{\text{effective}}^{(t)}=\Theta^{(t)}-\Theta^{(0)}$ on every optimizer state,161616 In the distributed case, this might not incur any additional cost, since $\Theta^{(0)}$ would already need to be stored. But this would require solving a linear system on every iteration. or one could store and continually update $\Lambda_{\text{effective}}$ as part of the optimizer state.
Appendix C Additional Results
Figure 2: Full results for gradient reconstruction error (i.e., $\|\overline{\Theta}-{\mathbf{S}}^{\top}{\mathbf{S}}\overline{\Theta}\|^{2}$) (left) and cosine similarity (i.e., $\cos(\overline{\Theta},{\mathbf{S}}^{\top}{\mathbf{S}}\overline{\Theta})$) of the various transformations across training steps with the 200M model. The projections with lowest reconstruction error (measured either by L2 error or cosine similarity with the unprojected stochastic gradient) do not give the best downstream performance (see Tab.˜2).
Appendix D Experimental Setup
The details of the two architectures we consider are shown in Tab.˜5. We also include a discussion on gradient accumulation.
@@ Appendix D Experimental Setup
The experimental setup in the original GaLore paper did not perform gradient accumulation, which meant that the maximum sequence length had to be short enough (e.g., 256) such that a single batch could contain a large-enough number of sequences for accurate gradient estimation. Our experiments are instead conducted in the standard setting where we assume gradients are accumulated across multiple microbatches. In this case, the reparameterization of GaLore as LoRA has the additional benefit of straightforwardly allowing for gradient accumulation in the lower-dimensional space. Concretely, the most straightforward implementation of GaLore [^17] E.g., the official implementation in https://github.com/jiaweizzhao/GaLore. will lead to gradient accumulation in the original parameter space, which would consume substantially more memory. In contrast, in the LoRA formulation the gradients are accumulated after applying the gradient transformation, providing substantial memory savings without any additional code. [^18] One can implement this optimization in the GaLore form, but this requires additional code. The issue is that most deep learning frameworks will compute the gradients of a parameterized function, and the user then separately passes these as input to an optimizer. If gradients are only transformed in the optimizer, then the automatic differentiation module cannot figure out that only the smaller transformed gradient is needed, and not the full gradient. We suspect that a sufficiently good compiler should in principle recover this optimization if one ensures that entire training steps (viz., all the gradient accumulation steps and optimizer step) are compiled jointly.